User Tools

Site Tools


lbaops:lbanov2012:v252ahcdlog

Agilent set to 17.4 GHz

SML01/02 set to 336 MHz

No coherence confirmed at either 892.2 or 929.375 MHz. (Problem later confirmed to be with LO signal generator used to inject tone.) No fringes confirmed in vt02r tests, but confirmed later on experiment source.

Recording to CURT V004A.

NTP time incorrect until 11:39:27 (UT00:39:27) (offset by -25 seconds) due to NTP server failure. Recording stopped, ntp resynched against fixed NTP server & recording resumed.

Halted schedule between UT 01:39 and UT 01:52 to test for presence of coherence tone. Failed to see any tone anywhere in the bandpass. Back on source by UT 01:57.

Windstowed from ~06:00 UT until ~07:10 UT.

Missed all data from UT 21:50 to UT 22:40 due to FS timing error.

Disk changed to ATNF V012A at ~ UT 01:20

RCP (Tsys 1) fluctuating bet. 2500 and 9000 Jy. This was noted during the changeover to the following experiment and it was likely to be fluctuating for the duration of this one as well.

lbaops/lbanov2012/v252ahcdlog.txt · Last modified: 2015/12/18 16:38 (external edit)